OpenStudy (anonymous):

solve by substitution method: 9x+7y=-24 x=19-8y what is the solution of systems

OpenStudy (anonymous):

substitute 19 - 8y on the variable x of the first equation. so, it will look like this: 9(19-8y) + 7y = -24 then, you can solve for y now.

OpenStudy (anonymous):

\[9(19-8y)+7y=-24\] etc

OpenStudy (anonymous):

what awesome one said. now you have one equation with one unknown

OpenStudy (anonymous):

I couldn't solve it

OpenStudy (anonymous):

ok multiply out using distributive law

OpenStudy (anonymous):

\[38-72y+7y=-24\] collect like terms \[38-65y=-24\]

OpenStudy (anonymous):

subtract 38 \[-65y=-62\] divide \[y=\frac{62}{65}\]

OpenStudy (anonymous):

i made a mistake

OpenStudy (anonymous):

\[ 9(19-8y) + 7y = -24 \]

OpenStudy (anonymous):

\[171-65y=-24\] \[-65y=-24-171=-195\] \[y=\frac{-195}{-65}=3\]

OpenStudy (anonymous):

now since y = 3 substitute back in equation 2 to get \[x=19-8\times 3=-5\]

OpenStudy (anonymous):

so satellite73 the solution ordered pair is (-5,3) correct

OpenStudy (anonymous):

so satellite73 the solution ordered pair is (-5,3) correct

OpenStudy (anonymous):

yes that is it

OpenStudy (anonymous):

thanks again